# Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching

### v1

@@ Abstract
Entity matching (EM) is a critical step in entity resolution. Recently, entity matching based on large language models (LLMs) has shown great promise. However, current LLM-based entity matching approaches typically follow a binary matching paradigm that ignores the global consistency between different records. In this paper, we investigate various methodologies for LLM-based entity matching that incorporate record interactions from different perspectives. Specifically, we comprehensively compare three representative strategies: matching, comparing, and selecting, and analyze their respective advantages and challenges in diverse scenarios. Based on our findings, we further design a compositional entity matching (ComEM) framework that leverages the composition of multiple strategies and LLMs. In this way, ComEM can benefit from the advantages of different sides and achieve improvements in both effectiveness and efficiency. Experimental results show that ComEM not only achieves significant performance gains on various datasets but also reduces the cost of LLM-based entity matching in real-world scenarios.
@@
Tianshu Wang1,2, Hongyu Lin1, Xiaoyang Chen3, Xianpei Han1,4, Hao Wang5, Zhenyu Zeng5, Le Sun1,4 1 Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2 Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences 3 University of Chinese Academy of Sciences 4 State Key Laboratory of Computer Science, Institute of Software, Chinese Academy of Sciences 5 Alibaba Cloud Group {tianshu2020, hongyu, xianpei, sunle}@iscas.ac.cn, chenxiaoyang19@mails.ucas.ac.cn, cashenry@126.com, zhenyu.zzy@alibaba-inc.com
@@ 1 Introduction
Figure 1: Three strategies for EM with LLMs.
@@
However, current LLM-based entity matching methods identify matches by classifying each pair of records independently. This matching strategy ignores the global consistency between records and thus leads to suboptimal results. On the one hand, ER applications require more than independent classification due to the interconnected nature of record relationships (Getoor and Machanavajjhala, 2012). For example, in record linkage (a.k.a. clean-clean ER), a single record from one data source typically matches at most one record from another data source, because there are usually no duplicates in a single database. Unfortunately, matching-based approaches do not take advantage of this nature of record linkage well. On the other hand, this strategy ignores the capabilities of LLMs to handle multiple records simultaneously to distinguish similar records. Using the records in Figure 1(c) as an example, if “Cruzer Glide”, “Cruzer Force”, and “Extreme Pro” appear in different records of the same context, LLMs are more likely to recognize that these are different models of SanDisk flash drives and use them as the basis for identifying matches. As a result, the matching strategy cannot fully unleash the potential of LLMs in EM.
@@
In this paper, we thoroughly investigate three strategies for entity matching with LLMs that incorporate record interactions from different perspectives, as shown in Figure 1. Specifically, apart from the conventional matching strategy shown in Figure 1(a), we investigate two additional strategies that leverage information from other records: 1) the comparing strategy, which identifies the record out of two that is more likely to match the anchor record, as shown in Figure 1(b); 2) the selecting strategy, which directly chooses the record from a list that is most likely to match the anchor record, as shown in Figure 1(c). Our research suggests that for entity matching with LLMs, incorporating record interactions is critical and can significantly improve EM performance in various scenarios. Therefore, the global selecting strategy is often the most effective. However, we also observe that the selecting strategy suffers from the position bias and limited long context understanding of current LLMs, which restricts its generality in applications.
@@
Based on our findings, we design a compositional entity matching (ComEM) framework that leverages the composition of multiple strategies and LLMs for both effectiveness and efficiency. Specifically, given an entity record and its $n$ candidate matches, we first preliminarily rank and filter these candidates using the local matching or comparing strategy, implemented with a medium-sized LLM. We then perform fine-grained identification on the top $k$ candidates using the global selecting strategy, facilitated by a powerful LLM. This approach not only mitigates the challenges and biases faced by the selecting strategy with too many options, but also reduces the cost of LLM invocations caused by composing multiple strategies. Consequently, by integrating the advantages of different strategies and LLMs, ComEM achieves a more effective and efficient entity matching process.
@@
To investigate different strategies with LLMs and to evaluate our ComEM framework, we conducted in-depth experiments on 8 ER datasets. Experimental results verify the effectiveness of incorporating record interactions through the selecting strategy, with an average 13.39% improvement in F1 over the current matching strategy. In addition, we examined the effect of 10 different LLMs using these strategies on identification and ranking. Based on the results, our ComEM framework is able to further improve the average F1 of the single selecting strategy by up to 8.46% while reducing the cost.
@@
Generally speaking, our contributions can be summarized as follows:
  • We investigate three strategies for entity matching with LLMs, and delve into their advantages and shortcomings in different scenarios.
@@ 2.1 Entity Resolution
As the core of data integration and cleaning, entity resolution has received extensive attention over the past six decades (Fellegi and Sunter, 1969; Getoor and Machanavajjhala, 2012; Binette and Steorts, 2020; Papadakis et al., 2021). The blocking-and-matching pipeline has become the mainstream of entity resolution, where blocking filters out obviously dissimilar records and matching identifies duplicates through complex techniques.
Blocking.
Traditional blocking approaches group records into blocks by shared signatures, followed by cleaning up unnecessary blocks and comparisons (Papadakis et al., 2022). Meta-blocking further reduces superfluous candidates by weighting candidate record pairs and graph pruning (Papadakis et al., 2014). Recently, nearest-neighbor search techniques, especially cardinality-based, have gained more attention and achieved the state-of-the-art results for blocking (Thirumuruganathan et al., 2021; Paulsen et al., 2023).
Entity Matching.
The open and complex nature of entity matching has spurred the development of a variety of approaches to address this persistent challenge, including rule-based (Benjelloun et al., 2009; Li et al., 2015), distance-based (Bilenko et al., 2003), and probabilistic methods (Fellegi and Sunter, 1969; Wu et al., 2020), etc. With the introduction of deep learning methods (Mudgal et al., 2018), especially pre-trained language models (Li et al., 2020), entity matching has made significant progress (Barlaug and Gulla, 2021; Tu et al., 2023). The emergence of LLMs brings a new zero- or few-shot paradigm to entity matching (Narayan et al., 2022; Peeters and Bizer, 2023a), alleviating training data requirements and the unseen entity problem (Wang et al., 2022; Peeters et al., 2024). Most deep learning and LLM-based approaches treat entity matching as an independent classification problem, except for GNEM (Chen et al., 2021) which models this task as a collective classification task on graphs. To the best of our knowledge, this is the first effort to formulate entity matching as a comparison or selection task with LLMs.
@@ 2.2 Large Language Model
The advent of LLMs like ChatGPT marks a significant advance in artificial intelligence, offering unprecedented natural language understanding and generation capabilities, and even general intelligence (Bubeck et al., 2023). By scaling up the model and data size of pre-trained language models, LLMs exhibit emergent abilities (Wei et al., 2022) and can thus solve a variety of complex tasks by “prompt engineering” without “fine-tuning”. For more technical details on LLMs, we refer the reader to the related survey (Zhao et al., 2023).
3 Problem Formulation
In this paper, we formulate the task of entity matching as the process of identifying matching records from a given entity record $r$ and its candidate matches $R_{c}=\{r_{1},r_{2},\ldots,r_{n}\}$. This formulation mitigates the limitations of independent pairwise matching and responds to real-world entity resolution scenarios. Current state-of-the-art blocking methods adhere to the k-nearest neighbor (kNN) search paradigm, which retrieves a list of candidate matches for each entity record by transforming it into a sparse or dense vector and performing the kNN search. In addition, this formulation accommodates both single-source deduplication and dual-source record linkage, and makes good use of the 1-1 assumption, i.e., record $r$ matches at most one of the candidate matches $R_{c}$. This assumption is widespread in record linkage, and deduplication with canonical entity construction.
@@ 4 Entity Matching with LLMs
In this section, we first introduce three strategies for entity matching with LLMs. We then propose our ComEM framework that leverages the composition of multiple strategies and LLMs.
@@ 4.1 LLM as a Matcher
Recent work formulates entity matching as a binary classification task based on LLMs (Narayan et al., 2022; Peeters and Bizer, 2023b, a; Fan et al., 2023; Li et al., 2024). In this strategy, an LLM acts as a pairwise matcher to determine whether two records match. Specifically, given an entity record $r$ and its candidate matches $R_{c}=\{r_{1},r_{2},\ldots,r_{n}\}$, this approach independently classifies each pair of records $(r,r_{i})_{1\leq i\leq n}$ as matching or not by interfacing LLMs with an appropriate matching prompt $\operatorname{P}_{m}$, as shown in Figure 1(a):
$\operatorname{LLM}_{m}\colon\left\{(r,r_{i})\mid r_{i}\in R_{c}\right\}\xrightarrow{\operatorname{P}_{m}}\left\{\mathrm{Yes},\mathrm{No}\right\}$
Unlike previous studies, the core of LLM-based applications is to prompt LLMs to generate the correct answer, namely prompt engineering. An appropriate prompt should include the task instruction, such as “Do these two records refer to the same real-world entity? Answer Yes or No.” Optionally, a prompt could include detailed rules or several in-context learning examples to guide LLMs in performing this task. Given the need for long contexts in other strategies, and the instability of existing prompt engineering methods for entity matching Peeters and Bizer (2023a), we only attempt few-shot prompting on the matching strategy and leave the exploration of better prompt engineering with different strategies to future work.
This independent matching strategy ignores the global consistency of ER, as well as the capabilities of LLMs to incorporate record interactions. For record linkage, according to the well-known 1-1 assumption Getoor and Machanavajjhala (2012), each entity record $r$ matches at most one record of the candidate matches $R_{c}$. For deduplication, this restriction of at most one match per record can also be satisfied by constructing canonical entities. The traditional solution to satisfy these constraints is to construct a graph based on the similarity scores $s_{i}$ of record pairs $(r,r_{i})$ and to further cluster on the similarity graph. We can obtain the similarity scores from LLMs by calibrating the generated probabilities $p$ of labels Qin et al. (2023). Formally, the similarity score $s_{i}$ can be defined as:
$s_{i}=\begin{cases}1+p(\mathrm{Yes}\mid\operatorname{P}_{m}(r,r_{i})),&\text{if generate ``Yes''}\\ 1-p(\mathrm{No}\mid\operatorname{P}_{m}(r,r_{i})),&\text{if generate ``No''}\end{cases}$
Unfortunately, the generation probabilities are not available for many black-box commercial LLMs. Moreover, the probabilities on short-form labels are misaligned for common open-source chat-tuned LLMs because they are fine-tuned to respond in detail. The need to investigate better strategies for entity matching with LLMs arises in ER.
@@ 4.2 LLM as a Comparator
In this section, we introduce a comparing strategy for LLM-based entity matching that simultaneously compares two candidate matches to a given record. Specifically, given an entity record $r$ and its candidate matches $R_{c}=\{r_{1},r_{2},\ldots,r_{n}\}$, the comparing strategy compares two records $r_{i}$ and $r_{j}$ from candidate matches $R_{c}$ to determine which is more consistent with record $r$ by interfacing LLMs with a comparison prompt $\operatorname{P}_{c}$, as shown in Figure 1(b):
$\operatorname{LLM}_{c}\colon\left\{(r,r_{i},r_{j})\mid r_{i,j}\in R_{c}\right\}\xrightarrow{\operatorname{P}_{c}}\left\{\mathrm{A},\mathrm{B}\right\}$
where $\mathrm{A}$ and $\mathrm{B}$ are the labels corresponding to record $r_{i}$ and $r_{j}$. Since LLMs are sensitive to the prompt order, we compare each pair of records $(r_{i},r_{j})$ to record $r$ twice by swapping their order, i.e., $\operatorname{LLM}_{c}(r,r_{i},r_{j})$ and $\operatorname{LLM}_{c}(r,r_{j},r_{i})$.
Compared to the matching strategy, this comparing strategy introduces an additional record for more record interactions and shifts the task paradigm. It focuses on indicating the relative relationship between two candidate matches of a given record, rather than making a direct match or no match decision. Therefore, this strategy is suitable for ranking and fine-grained filtering to determine the most likely records for identification.
@@
$s_{i}=\sum_{j\neq i}\left(p(\mathrm{A}\mid\operatorname{P}_{c}(r,r_{i},r_{j}))+p(\mathrm{B}\mid\operatorname{P}_{c}(r,r_{j},r_{i}))\right)$
Here, probabilities are incorporated to calculate similarity scores, integrating a more nuanced assessment from the comparisons made by LLMs.
However, the advantage of the comparing strategy in obtaining similarity scores comes at the cost of using LLMs as the basic unit of comparison and $\mathcal{O}(n^{2})$ complexity. Fortunately, for entity matching, we only care about a small number of most similar candidates, and there are many comparison sort algorithms available to find top-$k$ elements efficiently. In this paper, we use the bubble sort algorithm to find the top-$k$ elements, optimizing the complexity of the comparing strategy to $\mathcal{O}(kn)$. To avoid confusion, we refer to the comparison of all pairs as comparingall-pair in experiments.

### v2

@@ Abstract
Entity matching (EM) is a critical step in entity resolution (ER). Recently, entity matching based on large language models (LLMs) has shown great promise. However, current LLM-based entity matching approaches typically follow a binary matching paradigm that ignores the global consistency between record relationships. In this paper, we investigate various methodologies for LLM-based entity matching that incorporate record interactions from different perspectives. Specifically, we comprehensively compare three representative strategies: matching, comparing, and selecting, and analyze their respective advantages and challenges in diverse scenarios. Based on our findings, we further design a compound entity matching framework (ComEM) that leverages the composition of multiple strategies and LLMs. ComEM benefits from the advantages of different sides and achieves improvements in both effectiveness and efficiency. Experimental results on 8 ER datasets and 9 LLMs verify the superiority of incorporating record interactions through the selecting strategy, as well as the further cost-effectiveness brought by ComEM.
@@
Tianshu Wang1,2, Xiaoyang Chen3, Hongyu Lin1, Xuanang Chen1, Xianpei Han1,4, Hao Wang5, Zhenyu Zeng5, Le Sun1,4 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences 3University of Chinese Academy of Sciences 4State Key Laboratory of Computer Science, Institute of Software, Chinese Academy of Sciences 5Alibaba Cloud Intelligence Group {tianshu2020, hongyu, xuanang2020, xianpei, sunle}@iscas.ac.cn, chenxiaoyang19@mails.ucas.ac.cn, cashenry@126.com, zhenyu.zzy@alibaba-inc.com
@@ 1 Introduction
Figure 1: Three strategies for LLM-based entity matching. We omit other attributes of records for simplicity.
@@
However, current LLM-based entity matching methods identify matches by classifying each pair of records independently. This matching strategy ignores the global consistency between record relationships and thus leads to suboptimal results. On the one hand, entity resolution requires more than independent classification due to the interconnected nature of record relationships (Getoor and Machanavajjhala, 2012). For example, in record linkage (i.e., clean-clean ER), a single record from one data source typically matches at most one record from another data source, since there are usually no duplicates in a single database. Unfortunately, matching-based approaches do not take advantage of this nature of record linkage. On the other hand, this strategy ignores the capabilities of LLMs to handle multiple records simultaneously to distinguish similar records. Using the records in Figure 1(c) as an example, if “Cruzer Glide”, “Cruzer Force”, and “Extreme Pro” appear in different records of the same context, LLMs are more likely to recognize that they are different SanDisk flash drive models, which helps with accurate matching. As a result, the matching strategy cannot fully unleash the potential of LLMs in EM.
@@
In this paper, we thoroughly investigate three strategies for LLM-based entity matching that incorporate record interactions from different perspectives, as shown in Figure 1. Specifically, apart from the conventional matching strategy shown in Figure 1(a), we investigate two additional strategies that leverage information from other records: 1) the comparing strategy, which identifies the record out of two that is more likely to match the anchor record, as shown in Figure 1(b); 2) the selecting strategy, which directly chooses the record from a list that is most likely to match the anchor record, as shown in Figure 1(c). Our research suggests that for LLM-based entity matching, incorporating record interactions is critical and can significantly improve EM performance in various scenarios. Therefore, the global selecting strategy is often the most effective. Nevertheless, we also observe that the selection accuracy decreases greatly as the position of the matching record increases in the candidate list. The position bias and limited long context understanding of current LLMs hinder the generality of the selecting strategy.
@@
Based on our findings, we design a compound entity matching framework (ComEM) that leverages the composition of multiple strategies and LLMs. Specifically, given an entity record and its $n$ potential matches obtained from the blocking step, we first preliminarily rank and filter these candidates using the local matching or comparing strategy, implemented with a medium-sized LLM. We then perform fine-grained identification on only the top $k$ candidates using the global selecting strategy, facilitated by a more powerful LLM. This approach not only mitigates the challenges and biases faced by the selecting strategy with too many options, but also reduces the cost of LLM invocations caused by composing multiple strategies. Consequently, by integrating the advantages of different strategies and LLMs, ComEM achieves a more effective and efficient entity matching process.
@@
To investigate different strategies and to evaluate our ComEM framework, we conducted in-depth experiments on 8 ER datasets. Experimental results verify the effectiveness of incorporating record interactions through the selecting strategy, with an average 17.58% improvement in F1 over the current matching strategy. In addition, we examined the effect of 9 different LLMs using these strategies on identification or ranking. Based on the results, ComEM is able to further improve the average F1 of the single selecting strategy by up to 4.01% while reducing the cost.
@@
Generally speaking, our contributions can be summarized as follows [^1] The source code of this paper is available at: anonymous.4open.science/r/LLM4EM and supplementary material.:
  • We investigate three strategies for LLM-based entity matching, and delve into their advantages and shortcomings in different scenarios.
@@ 2.1 Entity Resolution
As the core of data integration and cleaning, entity resolution has received extensive attention over the past decades (Fellegi and Sunter, 1969; Getoor and Machanavajjhala, 2012; Binette and Steorts, 2020; Papadakis et al., 2021). The blocking-and-matching pipeline has become the mainstream of entity resolution, where blocking filters out obviously dissimilar records and matching identifies duplicates through complex techniques.
Blocking.  Traditional blocking approaches group records into blocks by shared signatures, followed by cleaning up unnecessary blocks and comparisons (Papadakis et al., 2022). Meta-blocking further reduces superfluous candidates by weighting potential record pairs and graph pruning (Papadakis et al., 2014). Recently, nearest-neighbor search techniques, especially cardinality-based ones, have gained more attention and achieved state-of-the-art (SOTA) results for blocking (Thirumuruganathan et al., 2021; Paulsen et al., 2023).
Entity Matching.  The open and complex nature of entity matching has spurred the development of various approaches to address this persistent challenge, including rule-based (Benjelloun et al., 2009; Li et al., 2015), distance-based (Bilenko et al., 2003), and probabilistic methods (Fellegi and Sunter, 1969; Wu et al., 2020), etc. With the advent of deep learning methods (Mudgal et al., 2018), especially pre-trained language models (PLMs) (Li et al., 2020), entity matching has made significant progress (Barlaug and Gulla, 2021; Tu et al., 2023). The emergence of LLMs brings a new zero- or few-shot paradigm to entity matching (Narayan et al., 2022; Peeters and Bizer, 2023a), alleviating training data requirements. Most deep learning and LLM-based approaches treat entity matching as an independent classification problem, except for GNEM (Chen et al., 2021), which models this task as a collective classification task on graphs. To the best of our knowledge, this is the first effort to formulate entity matching as a comparison or selection task using LLMs.
@@ 2.2 Large Language Model
The advent of LLMs such as ChatGPT marks a significant advance in artificial intelligence, offering unprecedented natural language understanding and generation capabilities, and even general intelligence (Bubeck et al., 2023). By scaling up the model and data size of PLMs, LLMs exhibit emergent abilities (Wei et al., 2022) and can thus solve a variety of complex tasks by “prompt engineering” without “fine-tuning”. For more technical details on LLMs, we refer the reader to the related survey (Zhao et al., 2023).
@@ 3 Entity Matching with LLMs
In this section, we first present the problem formulation. Then, we introduce three strategies for LLM-based entity matching. Finally, we propose our ComEM framework, which leverages the composition of multiple strategies and LLMs.
3.1 Problem Formulation
We formulate the task of entity matching as the process of identifying matching records from a given entity record $r$ and its $n$ potential matches $R=\{r_{1},r_{2},\ldots,r_{n}\}$ obtained from blocking. This formulation mitigates the limitations of independent pairwise matching and fits real-world entity resolution scenarios. First, current SOTA blocking methods adhere to the k-nearest neighbor (kNN) search paradigm, which retrieves a list of potential matches for each entity record, rather than generating candidate matches pairwise as in traditional blocking workflows. In addition, this formulation accommodates both single-source deduplication and dual-source record linkage, and makes good use of the 1-1 assumption, i.e., record $r$ matches at most one of the potential matches $R$. This assumption is widespread in record linkage, and deduplication with canonical entity construction.
@@ 3.2 LLM as a Matcher
Recent work formulates entity matching as a binary classification task based on LLMs (Narayan et al., 2022; Peeters and Bizer, 2023b, a; Fan et al., 2023; Li et al., 2024). In this strategy, an LLM acts as a pairwise matcher to determine whether two records match. Specifically, given an entity record $r$ and its potential matches $R=\{r_{1},r_{2},\ldots,r_{n}\}$, this approach independently classifies each pair of records $(r,r_{i})_{1\leq i\leq n}$ as matching or not by interfacing LLMs with an appropriate matching prompt, as shown in Figure 1(a):
$\operatorname{LLM}_{m}\colon\left\{(r,r_{i})\mid r_{i}\in R\right\}\rightarrow\left\{\mathrm{Yes},\mathrm{No}\right\}$
Unlike previous studies, the core of LLM-based applications is to prompt LLMs to generate the correct answer, namely prompt engineering. An appropriate prompt should include the task instruction, such as “Do these two records refer to the same real-world entity? Answer Yes or No”. Optionally, a prompt could include detailed rules or several in-context learning examples to guide LLMs in performing this task. Given the need for long contexts in other strategies, and the instability of existing prompt engineering methods for entity matching Peeters and Bizer (2023a), we only attempt few-shot prompting for the matching strategy and leave the exploration of better prompt engineering with different strategies to future work.
This independent matching strategy ignores the global consistency of ER, as well as the capabilities of LLMs to incorporate record interactions. For record linkage, according to the well-known 1-1 assumption, each entity record $r$ matches at most one record of the potential matches $R$. For deduplication, this assumption can also be satisfied by constructing canonical entities. The traditional solution to satisfy these constraints is to construct a graph based on the similarity scores $s_{i}$ of record pairs $(r,r_{i})$ and to further cluster on the similarity graph. We can obtain the similarity scores from LLMs by calibrating the generated probabilities $p$ of labels Qin et al. (2023). Formally, the similarity score $s_{i}$ can be defined as:
$s_{i}=\begin{cases}1+p(\mathrm{Yes}\mid(r,r_{i})),&\text{if generate ``Yes''}\\ 1-p(\mathrm{No}\mid(r,r_{i})),&\text{if generate ``No''}\end{cases}$
Unfortunately, the generation probabilities are not available for many black-box commercial LLMs. Moreover, the probabilities on short-form labels are misaligned for common open-source chat-tuned LLMs because they are fine-tuned to respond in detail. The need to investigate better strategies for LLM-based entity matching arises in ER.
@@ 3.3 LLM as a Comparator
In this section, we introduce a comparing strategy for LLM-based entity matching that simultaneously compares two potential matches to a given record. Specifically, given an entity record $r$ and its potential matches $R=\{r_{1},r_{2},\ldots,r_{n}\}$, the comparing strategy compares two records $r_{i}$ and $r_{j}$ from potential matches $R$ to determine which is more consistent with record $r$ by interfacing LLMs with a comparison prompt, as shown in Figure 1(b):
$\operatorname{LLM}_{c}\colon\left\{(r,r_{i},r_{j})\mid r_{i,j}\in R\right\}\rightarrow\left\{\mathrm{A},\mathrm{B}\right\}$
where $\mathrm{A}$ and $\mathrm{B}$ are labels corresponding to record $r_{i}$ and $r_{j}$. Since LLMs may be sensitive to the prompt order, we compare the record pair $(r_{i},r_{j})$ to record $r$ twice by swapping their order.
Compared to the matching strategy, the comparing strategy introduces an additional record for more record interactions and shifts the task paradigm. It focuses on indicating the relative relationship between two potential matches of a given record, rather than making a direct match or no match decision. Therefore, this strategy is suitable for ranking and fine-grained filtering to determine the most likely records for identification.
@@
$s_{i}=\sum_{j\neq i}\left(p(\mathrm{A}\mid(r,r_{i},r_{j}))+p(\mathrm{B}\mid(r,r_{j},r_{i}))\right)$
However, the advantage of the comparing strategy in obtaining similarity scores comes at the cost of using LLMs as the basic unit of comparison and $\mathcal{O}(n^{2})$ complexity. Fortunately, for entity matching, we only care about a small number of most similar candidates, and there are many comparison sort algorithms available to find the top-$k$ elements efficiently. In this paper, we use the bubble sort algorithm to find the top-$k$ elements, optimizing the complexity of the comparing strategy to $\mathcal{O}(kn)$. To avoid confusion, we refer to the comparison of all pairs as comparingall-pair in our experiments.